Shift Summary: 12/14 Owl

Jason LaDue (jladue@cuhep.Colorado.EDU)
Tue, 14 Dec 1999 06:05:58 -0700 (MST)


Shift Summary for owl shift, 12/14/99

On Shift: U. Erkarslan, J. LaDue

Hours of beam available at all intensities: 4.00
Hours of beam available at nominal intensity: 3.50
Hours of E799 data (special runs not included): 3.75
Hours of special runs (laser scans, muon runs...): 0.00

Accounting of lost time:
Accelerator problems with Booster power supplies accounts for
most of the time. On our end is the usual assortment of KFRENDS and
pipeline inits.

========================================================================
Runs ended this shift:
15216 - X.XXXM - BEAM_799 -> no actual beam during this run.
15217 - 0.430M - BEAM_799 -> stopped while accel down to change tapes

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:
None

========================================================================
Overview of shift:

First half of shift is plagued with Booster problems, which
prevent any beam. At the beginning of second half of shift the intensity
is quite high (1.05E13), but then it goes down to nominal. Mostly
quiet after that, though we are still getting the occasional scaler
alarms.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
00:00 No Beam due to Booster RF problems

15217 01:28 Run started and Beam comes back at Low (4E12)
intensity.
15217 01:57 Booster Power problems, no beam.

02:35 Suddenly realizing that he is in charge
Jason, drunk with power, stops the run and
changes tapes (they are ~72% full).

15218 03:23 Beam returns, intensity is high(~1E13).

04:26 First notice problems with CsI 961 comparator.
It is on the IDIGNORE list so I send email to
experts.

05:00 Split Tape UPF056 mounted.

83-4 05:01 No Triggers, HCC buffer is full. KFREND and
continue.

05:36 Notice funny beam intensity Structure. SEM is
changing from spill to spill in a periodic man-
ner. IE one spill SEM is 9.1 next is 9.5, repeat.
MCR is notified.

05:45 RC8 and SA4 scaler alarms still go off occasion
ally

========================================================================
Follow up Notes on Previous Shift
CsI reference plots updated, will take affect next run.
Data taken.

========================================================================
Items to Follow up on Next Shift
CsI 961 has comparator problems, keep an eye on it.
SA3 may need its constants updated.